How do I calculate inpedance in a circuit


1
Expert's answer
2022-01-21T08:25:55-0500

Impedance is given as follows:


"\hat Z(j \omega)\ = R + j \left(- \frac {1}{\omega C} + \omega L \right)"

where "R, C, L" are the total resistance, capacitance and inductance of the circle respectively, and "\omega" is the frequency of the signal.